Grant thinks of an odd number between 48 and 64. the tens digit is greater than the ones digit and the difference between the two digits is 4. What is Grant's Number?
Answer by ankor@dixie-net.com(22740)

let t = tens digit
let u = unit
:
an odd number between 48 and 64.
t = 4, 5 or 6
the tens digit is greater than the ones digit
t > u; 49 is excluded
t = 5 or 6
and the difference between the two digits is 4.
t - u = 4
t = 6
6 - u = 4
u = 2, then 62, but that is an even number
t = 5
5 - u = 4
u = 1, then 51 has to be the number